1. Calculate the mass of potassium chloride produced if 4.00 grams of KClO3 are completely decomposed according to the reaction: 2 KClO3   🡪  2KCl  +  3O2
1
Expert's answer
2021-01-20T03:16:34-0500

nKClO3=nKCl; "{\frac {m(KClO3)} {M(KClO3)}=\frac {m(KCl)} {M(KCl)}}" "m(KCl)=\frac {m(KClO3) ∗ M(KCl)} {M(KClO3)}" ; m(KCl) = 2.433 (grams).


SOLUTION: 2.433 grams